CARS MICROSCOPY: COMING OF AGE

\underline{X. SUNNEY XIE}, Harvard University, Department of Chemistry and Chemical Biology, Cambridge, MA 02138.

Coherent anti-Stokes Raman scattering (CARS) microscopy is an imaging technique based on nonlinear Raman spectroscopy. The vibrational contrast, high sensitivity, and three dimensional resolution make CARS microscopy a powerful tool for imaging live cells and tissues. Over the past few years, CARS microscopy has evolved into a matured technology with its ever-growing applications both in cell biology and biomedical research.